The electrical flux density between two plates separated by polystyrene of relative permittivity 2.5 is 5µ C/m2 . Find the voltage gradient between the plates. 


1
Expert's answer
2021-03-02T18:06:01-0500

The voltage gradient is:


"E=\frac{D}{\varepsilon_0 \varepsilon},"


where D - the electrical flux density;

"\varepsilon_0" - the vacuum permittivity;

"\varepsilon" - relative permittivity of polystyrene.


"E=\frac{5\cdot 10^{-6}}{8.85\cdot 10^{-12} \cdot 2.5}\approx 226\space kV/m."


Answer: 226 kV/m